The beer distributor also sells cases of Brew With The Unnecessarily Pretentious Long Name beer (“BWTUPLN” for short), made by a small regional brewery. The demand for BWTUPLN is normally distributed with a mean of 12 cases and a standard deviation of 4 cases per day. When placing a replenishment order (directly with the brewer), it takes an average of 4 days for the order to arrive, with a standard deviation of 2 days (assume the delivery time is normally distributed). The distributor uses an (s,Q) policy to manage its inventory of the cases of beer.

a) If the distributor desires a 97% service level, what should the reorder point be?

b) If the standard deviation of the lead-time was 1 day (instead of 2), what would the reorder point be?

c) If the lead-time was exactly 4 days, what would the reorder point be?
Answer #1

Q1:

Parameters

Quantity

Cycle service Level

CSL

0.97

Z-value

z

=+NORM.S.INV(0.97)

1.88

Lead time

L (days)

4.00

Days

Standard Deviation of lead time

σL

2

days

Average demand during day

d

12

0

Standard deviation of demand

σd

4

gallons per week

Continuous review model

Safety Stock

SS = z√(Lσd2 + dσL2)

1.88*√(4*42 + 12*22)

= 19.89

ROP

R = d*L + SS

12*4 + 19.89

= 67.89

Q2:

Parameters

Quantity

Cycle service Level

CSL

0.97

Z-value

z

=+NORM.S.INV(0.97)

1.88

Lead time

L (days)

4.00

Days

Standard Deviation of lead time

σL

1

days

Average demand during day

D

12

0

Standard deviation of demand

σd

4

gallons per week

Continuous review model

Safety Stock

SS = z√(Lσd2 + dσL2)

1.88*√(4*42 + 12*12)

= 16.389

ROP

R = d*L + SS

12*4 + 10.58

= 64.389

Q3:

Parameters

Quantity

Cycle service Level

CSL

0.97

Z-value

z

=+NORM.S.INV(0.97)

1.88

Lead time

L (days)

4.00

Days

Standard Deviation of lead time

σL

0

days

Average demand during day

D

12

0

Standard deviation of demand

σd

4

gallons per week

Continuous review model

Safety Stock

SS = z√(Lσd2 + dσL2)

1.88*√(4*42 + 12*02)

= 14.4

ROP

R = d*L + SS

12*4 + 14.4

= 62.4